The length of a rectangle is 3 feet less than its width. It's perimeter is 98 feet. Find the dimensions of the rectangle.
Answer by tanshar(1)

l=w-3
2(l+w)=98
2(2w-3)=98
2w-3=49
2w=52
w=26 feet
l=23 feet